2. On January 1, Clayton Cranes purchased a crane for $190,000. Clayton expects the crane to remain useful for ten years (1,22. On January 1, Clayton Cranes purchased a crane for $190,000. Clayton expects the crane to remain useful for ten years (1,2is this correct? please help with part c.
Answer #1

As Part a and b are correct. I have only done Part c

c)

( Net book value - Residual value ) * 2*(1/10) = Double-declining-balance depreciation
Yr1 ( 190000 - 0 ) * 2*(1/10) = 38000
Yr2 ( 152000 - 0 ) * 2*(1/10) = 30400

Year 1

Depreciation= $190000*2*(1/10)

= $190000*2/10= $38000

Year 2

Net book value= $190000-38000= $152000

Depreciation= $152000*2*(1/10)

= $152000*2/10= $30400